scenario 1 for questions 11 and 12

A monopolist faces the following demand curve, marginal revenue curve and total cost curve for its product:

Q = 400 - 2P

MR = 200 - Q

TC = 10Q


11. Refer to scenario 1. What is the profit-maximising level of output?

[1] 0

[2] 10

[3] 190

[4] 200

[5] 400



12. Refer to scenario 1. How much profit does the monopolist earn?

[1] R1 900,00

[2] R19 987,50

[3] R18 050,00

[4] R19 950,00


1
Expert's answer
2021-05-20T11:32:47-0400

Given

Q = 400 - 2P

MR = 200 - Q

TC = 10Q

"MC = \frac{dTC}{dQ} = 10"



(11)

Profit maximizing level of output is where MR=MC

200-Q = 10

Q=190 units

Answer = 3) 190


(12)

Profit earned

"Profit=(P\times Q)-(TC)"

Substitute the value of Q in Q = 400 - 2P to get price

"190=400-2P\\2P=400-190\\2P=210\\P=105\\profit=(P\times Q)-(TC)\\(105\times 190)-(10\times 190)\\=19,950-1,900\\=18,050\\"

Answer=[3] R18,050.00